OpenStudy (anonymous):

Using the Remainder Theorem. Determine if (x + 1) is the factor of f(x) = x^2 + 11x + 18

OpenStudy (loser66):

8

OpenStudy (anonymous):

how did you get that?

OpenStudy (loser66):

plug -1 into the function, if the answer is 0 so, -1 is the root, if not the answer is the remainder

OpenStudy (loser66):

that means (x+1) is not factor of the function.

OpenStudy (anonymous):

where did you get -1 from?

OpenStudy (loser66):

from (x+1)

OpenStudy (loser66):

got it?

OpenStudy (anonymous):

so when you plugged 8 into the function you got 0 so 8 is the answer?

OpenStudy (loser66):

nope. you misunderstand the concept. if -1 is a root, then x-(-1) is a factor of the function. At that time, function divided by (x-(-1)) =0; backward, if the division has the remainder, that means -1 is not a root, and then (x-(-1) is not a factor of thw function. that's why they ask you to use remainder theorem to consider whether it's a factor or not

OpenStudy (anonymous):

oh okay, so (-1)^2 + 11(-1) + 18 =8?

OpenStudy (loser66):

that means (x+1)|x^2 +11x+18 = 8. or, \[x^2 +11x+18=something +\frac{8}{x+1}\]

OpenStudy (loser66):

the remainder is 8

OpenStudy (anonymous):

so (x+1) is not a factor because the remainder is not 0?

OpenStudy (loser66):

if you plug the number into the function and get 0, so that (x-number) is factor of function

OpenStudy (loser66):

yup yes yeah

OpenStudy (anonymous):

oh okay, I understand now! Thank you so much!

OpenStudy (loser66):

yw